OpenStudy (anonymous):

find three consecutive odd integers whose sum is 363

OpenStudy (anonymous):

119+121+123=363

OpenStudy (anonymous):

how did you get those numbers

OpenStudy (anonymous):

Suppose that k is our odd number. Then k + 2 is the next odd number. And k + 4 is the third odd number. We want the sum of these numbers to be 363. So we want (k) + (k+2) + (k+4) = 363. So we want 3k + 6 = 363. This gives us 3k = 357, which gives k = 119.